Shift Summary: EVE, June 27, 1999

Alexander Ledovskoy (ledovsko@fnpxsr)
Mon, 28 Jun 1999 00:00:04 -0500

Shift Summary for EVE shift, 27-JUNE-99

On Shift: Andrea Soddu, Sasha Ledovskoy

Hours of beam delivered at all intensities: 8.5h
Hours of beam delivered at nominal intensity: 8.5h
Hours of E832 data (special runs not included): 7.5h
Hours of special runs (laser scans, muon runs...): 0.25h

Accounting of lost time:

Tape swap, ramping magnet, run inits

========================================================================
Runs ended this shift:

13745 BEAM_832 25.2M events
13746 KTEV_PED
13748 BEAM_MU_4E12
13749 BEAM_832 0.73M events

Tapes brought to FCC this shift:

QKC049-051
QKS049-051
QKK049-051
QKB017
QKA001
QKCC01-02
QKLA03-04
QKLS01
QKMU05-07
QKN001-004
QKNL01-04
QKPE02-03
QKSM01-02

========================================================================
Accesses during shift:

========================================================================
Overview of shift:

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

13745 427 3:30 Our shift crew takes over.
Shift starts with ONLINE SPLIT request to mount
QKCC02 on tps3d2. We took out QKSM02 from this
tape drive and successfully mounted the tape.

438 3:45 CSI 1887 has Laser stab. RMS 28.6%
Back to normal next spill.

466 4:25 ktevana errors. Missing FFFF from crate 53 stream 5
KFREND
Same ktevana errors on next spill
KFREND
Everything back to normal

507 HCC is inefficient error
Reload HCC+ETOT.
Back to normal.

525 HCC is inefficient error
Reload ETOT. Quick reset of the pipeline also
Back to normal.

540 HCC is inefficient error
Reload ETOT+HCC. Quick reset of the pipeline also
Back to normal. Sydney is notified about these
errors

568 HCC is inefficient error
Reload ETO
Back to normal.

7:05 Tapes are full. Run stops.

13746 Pedestal Run. Constants are updated for FERA ADCs

13748 Muon run

13749 7:55 Magnet polarity is changed to POS
BEAM_832 run started
Quite a few problems:
HCC is inefficient error
Pipeline errors
Crate 53 busy
KFRENDs don't help with crate 53.
Cr53 is cycled. Everything is OK now...
...But spill number hasn't changed
Stop this run and start new one

13750 8:37 BEAM_832
4 HCC is inefficient error - Reload Etot
7 CSI 907 RMS is 33.5% - Back to normal next spill

116 CSI 2415 rms is 31.4%
Back to normal next spill

122 CSI 967 rms is 31.1%
Back to normal next spill

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift